Marsha takes out a loan for 20,000. She is to repay the loan by level monthly payments of 1500 with the final payment possibly being less than the previous payments. The nominal interest rate is 12% convertible monthly. Let n be the number of payments and X be the amount of the final payment. (a) Determine n. (b) Determine X
Answer #1

Using financial calculator
PV=-20000
PMT=1500
I/Y=12%/12
FV=0
CPT N=14.38151613

So, total number of payments=15
X=20000*(1+12%/12)^15-1500/(12%/12)*((1+12%/12)^14-1)*(1+12%/12)=574.0358019